Quantitative Aptitude प्रश्न और उत्तर का अभ्यास करें
8प्र: If 25% of a number is subtracted from a second number, the second number reduces to its four-sixth. What is the ratio of the first number to the second number ? 2037 05b5cc6dee4d2b4197774ea46
5b5cc6dee4d2b4197774ea46- 14:3true
- 22:3false
- 33:2false
- 45:3false
- उत्तर देखेंउत्तर छिपाएं
- Workspace
- SingleChoice
उत्तर : 1. "4:3"
व्याख्या :
Answer: A) 4:3 Explanation: Let the numbers be p & q q-p/4 = 4q/6 => q-4q/6 = p/4 => 2q/6 = p/4 => p/q = 4/3
प्र: निम्न समीकरण को हल किजिए।
2036 05d4a95c9822faf1e4e7fd6ad
5d4a95c9822faf1e4e7fd6ad
- 11512false
- 21521true
- 31251false
- 41531false
- 5इनमें से कोई नहींfalse
- उत्तर देखेंउत्तर छिपाएं
- Workspace
- SingleChoice
उत्तर : 2. "1521"
प्र: 5 + 2 x 10/2 = ? 2034 05b5cc6a6e4d2b4197774cd33
5b5cc6a6e4d2b4197774cd33- 115true
- 235false
- 310false
- 420false
- उत्तर देखेंउत्तर छिपाएं
- Workspace
- SingleChoice
उत्तर : 1. "15"
व्याख्या :
Answer: A) 15 Explanation: Using BODMAS Rule, First solve division, mutiplication and Addition So here 10/2 = 5 Then 2 x 5 = 10 Now, 5 + 10 = 15. Hence, 5 + 2 x 10/2 = 15.
प्र: A man sell 35 waffers for Rs. 15 each and loses Rs. 40.4. If he sells 35 waffers for Rs. 18, how much does he loss or gain ? 2033 15b5cc74ae4d2b4197774fa95
5b5cc74ae4d2b4197774fa95- 1Rs. 64true
- 2Rs. 52false
- 3Rs. 72false
- 4Rs. 55false
- उत्तर देखेंउत्तर छिपाएं
- Workspace
- SingleChoice
उत्तर : 1. "Rs. 64"
व्याख्या :
Answer: A) Rs. 64 Explanation: Given Loss = Rs. 40.4We know loss = cost price - selling price 40.4 = CP - 35(15)=> CP = 40.4 + 525=> CP = Rs. 565.4Now the New selling price of waffer is Rs. 18=> SP = 35(18) = 630Now Gain = SP - CP = 630 -565.4 = Rs. 64.6.
प्र: A person walking 6/7 of his usual rate is 30 minutes late. What is his usual time ? 2031 05b5cc720e4d2b4197774f62c
5b5cc720e4d2b4197774f62c- 12 hrs 50 minfalse
- 23 hrs 25 minfalse
- 33 hrstrue
- 42 hrs 5 minfalse
- उत्तर देखेंउत्तर छिपाएं
- Workspace
- SingleChoice
उत्तर : 3. "3 hrs"
व्याख्या :
Answer: C) 3 hrs Explanation: Let t be the usual time. If the usual rate is 1 unit, then the unusual rate is 6/7.Balance by the distance,1 x t = (6/7)(t + 30)7t = 6(t + 30)we get t as,t = 180 minutes = 3 hrs.
प्र: 2500473+ (56)2÷7=? 2030 05b5cc7c8e4d2b41977750e36
5b5cc7c8e4d2b41977750e36- 1547false
- 2475false
- 3455false
- 4None of thesetrue
- उत्तर देखेंउत्तर छिपाएं
- Workspace
- SingleChoice
उत्तर : 4. "None of these"
व्याख्या :
Answer: D) None of these Explanation:
प्र: In a box, there are four marbles of white color and five marbles of black color. Two marbles are chosen randomly. What is the probability that both are of the same color? 2027 05b5cc6a3e4d2b4197774cc15
5b5cc6a3e4d2b4197774cc15- 12/9false
- 25/9false
- 34/9true
- 40false
- उत्तर देखेंउत्तर छिपाएं
- Workspace
- SingleChoice
उत्तर : 3. "4/9"
व्याख्या :
Answer: C) 4/9 Explanation: Number of white marbles = 4 Number of Black marbles = 5 Total number of marbles = 9 Number of ways, two marbles picked randomly = 9C2 Now, the required probability of picked marbles are to be of same color = 4C2/9C2 + 5C2/9C2 = 1/6 + 5/18 = 4/9.
प्र: A shopkeeper who deals in books sold a book at 16% loss. Had she charged an additional Rs.60 while selling it , her profit would have been 14%. Find the cost price, in rupees, of the book ? 2026 05b5cc723e4d2b4197774f677
5b5cc723e4d2b4197774f677- 1Rs. 185false
- 2Rs. 154false
- 3Rs. 200true
- 4Rs. 177false
- उत्तर देखेंउत्तर छिपाएं
- Workspace
- SingleChoice
उत्तर : 3. "Rs. 200"
व्याख्या :
Answer: C) Rs. 200 Explanation: Let the C.P be 'x'Then, the selling price S.P = x - 16x/100= 84x/100 = 21x/25Now, if the S.P is 60 more, then the profit is 14% => 21x/25 + 60 = x + 14x/100=> 114x/100 - 21x/25 = 60=> (57 - 42)x/50 = 60=> 15x/50 = 60x = 3000/15 = 200 Therefore, the Cost price C.P = x = Rs. 200

